The co-ordinates of a particle moving in xy-plane vary with time t as: x=at, y=bt+c. The curve traced by the particle is ---


a) Straight line


b) Parabola


c) Hyperbola


d) Circle

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Abhingya Patra

x=at ..........1

y=bt+c..........2

put value of t from equation 1

 y = b(x/a)  +c

 y =(b/a) x +c

 this is equation of straight line

so the traced curve is straight line

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin